### v1

# RefineStat: Efficient Exploration for Probabilistic Program Synthesis

Abstract
Probabilistic programming offers a powerful framework for modeling uncertainty, yet statistical model discovery in this domain entails navigating an immense search space under strict domain-specific constraints. When small language models are tasked with generating probabilistic programs, they frequently produce outputs that suffer from both syntactic, and semantic errors, such as flawed inference constructs. Motivated by probabilistic programmers’ domain expertise and debugging strategies, we introduce RefineStat, a language model–driven framework that enforces semantic constraints ensuring synthesized programs contain valid distributions, well-formed parameters, and then applies diagnostic-aware refinement by resampling prior or likelihood components whenever reliability checks fail. We evaluate RefineStat on multiple probabilistic-programming code-generation tasks using smaller language models (SLMs) and find that it produces programs that are both syntactically sound and statistically reliable, often matching or surpassing those from closed-source large language models (e.g., OpenAI o3).
1 Introduction
Scientific discovery often depends on our ability to express complex systems as statistical models. Finding appropriate models that are both interpretable and computationally efficient is challenging. The vision of automating model discovery has a long-standing history. Systems developed in the past have demonstrated success across various domains, such as identifying physical laws Bongard and Lipson (2007); McKinney et al. (2006); Linka et al. (2023), recovering the structure of nonlinear dynamical systems Schmidt and Lipson (2009), performing structure-aware nonparametric regression Duvenaud et al. (2013), and tackling unsupervised learning problems Grosse (2014). Despite these achievements, they typically relied on significant manual effort—experts were required to define a domain-specific language (DSL) for representing models and to engineer custom search algorithms for exploring compositions within that DSL.
Large Language Models (LLMs) have the potential to automate the model discovery by leveraging their extensive knowledge across various domains, enabling them to propose modeling approaches that were traditionally developed by human experts. However, using LLMs comes with significant challenges. Directly querying LLMs to generate statistical models often produces semantically flawed and unreliable programs, particularly in probabilistic programming languages like PyMC and NumPyro that evolve rapidly. These bugs hinder the correct execution of programs and constrain the effective exploration of the search space of working solutions. Further, running LLMs is costly, as expressive models (e.g., GPT-4) incur high API fees.
These limitations motivate Small Language Models (SLMs) as a practical alternative. Despite recent gains on coding tasks, they still produce semantic bugs code that runs but violates the intended statistical meaning. For example, in PyMC Salvatier et al. (2016), an SLM can generate code that places variance where a standard deviation is expected – pm.Normal(..., sigma=sigma**2) instead of pm.Normal(..., sigma=sigma). This change encodes the wrong statistical model, often inflating uncertainty and even triggering errors such as SamplingError. These issues motivate our constraints and Bayesian-workflow checks to detect and correct such errors.
Our Work: RefineStat We present RefineStat, a novel probabilistic programming synthesis framework that efficiently guides a LLM to generate probabilistic programs. RefineStat is the first to demonstrate that open-weight SLMs can synthesize reliable probabilistic programs in the Bayesian-workflow sense i.e., they satisfy standard checks. The checks include low split-$\widehat{R}$ and adequate bulk/tail ESS (good MCMC mixing), low NUTS divergences (well-behaved geometry), and strong out-of-sample fit via PSIS–LOO. These diagnostics plug into the loop of the Bayesian workflow (Gelman et al., 2020), and we adopt the improved $\widehat{R}$/ESS definitions (Vehtari et al., 2021) and PSIS–LOO for predictive assessment (Vehtari et al., 2017). Section 2 gives relevant definitions.
RefineStat produces reliable statistical programs through a two-phase approach: (1) semantically constrained generation and (2) diagnostic-aware refinement (Section 3). Our semantic constraining ensures that synthesized probabilistic programs contain valid distributions with well-formed parameters and proper variable dependencies. The diagnostic-aware refinement systematically resamples prior specifications or likelihood models when generated programs fail to meet established reliability criteria within the Bayesian workflow, thereby ensuring efficient search of probabilistic models using small langauge models.
We evaluate RefineStat on a suite of five representative probabilistic datasets, and five open-weight LLMs, with up to 8 billion weights. Our comparison shows that RefineStat significantly improves over directly querying LLMs in an unconstrained manner or only syntactic constraining with Syncode Ugare et al. (2024c). We show that the programs generated by RefineStat often pass the diagnostic metrics that indicate high quality to represent and explain the data. We also show that the RefineStat’s performance is comparable to a recent LLM-based generation algorithm BoxLM Li et al. (2024), which uses two GPT-4 LLM instances to iteratively propose a likely program and refine it, respectively; yet RefineStat obtains those results with a single small language model.
Contributions: The main contributions of this paper are:
  • Approach: We present RefineStat, a novel SLM-based framework for synthesis of probabilistic programs that are semantically correct and have high predictive performance.
  • Constrained decoding: We propose using semantic constrained decoding to help generate syntactically and semantically valid probabilistic programs, at a small overall cost.
  • Iterative program search: We present an iterative refinement loop that leverages a single, unmodified open-weights SLM to generate probabilistic programs with improved diagnostic metrics, refining statistical reliability by selectively resampling the likelihood and prior.
  • Evaluation: We demonstrate that RefineStat performs significantly better than baseline language models, in terms of different diagnostic metrics, and in some cases performs equally well as GPT-4 and hand-written developer programs.
2 Background
2.1 Language Models
Current autoregressive language models (LMs) operate on a vocabulary $V\subseteq\Sigma^{*}$ of tokens. A tokenizer converts an input prompt $O_{0}\in\Sigma^{*}$ into a sequence of tokens $t_{1},t_{2},\dots,t_{k}$. The LM $M:V^{*}\to\mathbb{R}^{|V|}$ takes this sequence and outputs scores $\mathcal{S}$ over the vocabulary: $\mathcal{S}=M(t_{1},t_{2},\dots,t_{k})$. A softmax function transforms these scores into a probability distribution, from which $t_{k+1}$ is sampled.
Decoding and Constraints. Various approaches for token selection include greedy decoding, sampling, and beam search, repeated until an end-of-sequence (EOS) token or another stopping criterion is met. In constrained decoding, we may need to exclude specific tokens at certain positions. This is achieved using a mask $m\in\{0,1\}^{|V|}$, where 1 indicates a viable token and 0 a discarded one. Decoding methods can then be applied to $m\odot\textit{softmax}(\mathcal{S})$.
Grammar-guided Generation Grammar-guided generation constrains model outputs to a formal grammar by using production rules of the form $A\to\alpha$, where $A$ is a nonterminal symbol and $\alpha$ is a sequence of nonterminals and terminals (the actual tokens or characters that appear in the final output). Most programming languages can be described using context-free grammar, with rules that apply to nonterminal symbols independent of their context. Grammar-guided generation ensures that LM outputs follow the syntactic structure required for, e.g., code generation or structured data formatting.
Bayesian Workflow A robust Bayesian analysis follows an iterative workflow of model formulation, posterior inference, diagnostic checking, and model comparison (Gelman et al., 2020). At its core, a generative program often follows a $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$ structure—first fixing the observed data ($\mathcal{D}$), then sampling latent variables under the prior ($\mathcal{P}:p(z)$), and finally conditioning on the data via the likelihood ($\mathcal{L}:p(x\mid z)$). Modern probabilistic programming languages such as Stan (Carpenter et al., 2017b) and PyMC (Salvatier et al., 2016) streamline this cycle by automating MCMC sampling and offering integrated diagnostics to assess prior implications, convergence (e.g., split-$\widehat{R}$, effective sample size, BFMI, divergent NUTS transitions) (Vehtari et al., 2021; Hoffman et al., 2014; Betancourt, 2017), and predictive accuracy via PSIS-LOO (Vehtari et al., 2017). These checks guard against mis-specification and poor sampling before models are deployed for inference or decision-making.
2.2 Probabilistic Programming
Statistical modeling aims to describe relationships between variables in data through joint probability distributions that capture both observed phenomena and underlying latent structure. In probabilistic modeling, we formalize this as a joint distribution $p(x,z|\eta)$, where $x=x_{1:n}$ represents $n$ observed data points, $z=z_{1:m}$ denotes $m$ latent variables, and $\eta$ corresponds to fixed model parameters. The inferential goal is to compute the posterior distribution $p(z|x)$, which quantifies uncertainty in the latent variables conditional on observed data. Probabilistic programming languages (PPL) provide a flexible computational substrate for specifying joint distributions $p(x,z\mid\eta)$ as programs while leveraging automated inference methods (e.g., MCMC, variational inference) to compute the posterior $p(z\mid x)$ van de Meent et al. (2021). They can be categorized as internal Domain-Specific Languages (DSLs), which embed within a host language and reuse its syntax and tooling (e.g., PyMC3 Salvatier et al. (2016), NumPyro Phan et al. (2019), Pyro Bingham et al. (2019)), or as external DSLs that define their own syntax and compiler (e.g., Stan Carpenter et al. (2017b)).
This representation is particularly valuable because it enables automated model specification while leveraging established inference algorithms Gordon et al. (2014).
Probabilistic Programming Metrics We briefly define the standard probabilistic programming diagnostics and metrics used in our framework. Detailed formal definitions are provided in Appendix A.
  1. 1.
    $\widehat{R}_{\phi}$: Split-$\widehat{R}$ statistic for parameter $\phi$, measuring MCMC chain convergence.
  2. 2.
    $\mathrm{ESS}_{\mathrm{bulk},\phi}$: The effective bulk sample size for the parameter $\phi$, estimating the sampling efficiency across the central mass of the posterior.
  3. 3.
    $\mathrm{ESS}_{\mathrm{tail},\phi}$: Tail effective sample size for parameter $\phi$, measuring sampling efficiency in tails.
  4. 4.
    $\mathrm{Divergences}(M)$: Count of divergent NUTS Hoffman et al. (2014) transitions in model $M$.
  5. 5.
    $\mathrm{BFMI}(M)$: Bayesian Fraction of Missing Information for model $M$, assessing energy transition efficiency in Hamiltonian Monte Carlo (HMC) Neal et al. (2011). HMC is a gradient-based MCMC algorithm that leverages Hamiltonian dynamics to propose coherent moves in parameter space, reducing random-walk behavior.
  6. 6.
    $\widehat{k}_{i}(M)$: Pareto shape parameter for observation $i$ in PSIS-LOO (Pareto-smoothed importance sampling leave-one-out cross-validation), quantifying reliability of importance sampling estimates. PSIS-LOO approximates exact LOO predictive densities by smoothing raw importance weights with a generalized Pareto fit to stabilize high-variance weights.
  7. 7.
    $\widehat{\mathrm{elpd}}$: Expected Log Pointwise Predictive Density under Leave-One-Out cross-validation, measuring model’s out-of-sample predictive accuracy.
To evaluate out-of-sample predictive accuracy, we rely on the expected log pointwise predictive density under leave-one-out cross-validation (ELPD-LOO). A direct computation of LOO requires refitting the model $n$ times (once for each observation), which is often too costly in practice. To avoid this, we use Pareto-smoothed importance sampling leave-one-out cross-validation (PSIS-LOO) (Vehtari et al., 2017), which provides a fast approximation to ELPD-LOO and also gives diagnostics on influential observations via the Pareto $\widehat{k}$ values:
Definition 1 (PSIS-LOO)
Let $\mathcal{D}=\{y_{i}\}_{i=1}^{n}$ be the observed data and $M$ a model. Draw $\{\theta^{(s)}\}_{s=1}^{S}\sim p(\theta\mid\mathcal{D})$, compute raw weights $w_{i}^{(s)}=1/p(y_{i}\mid\theta^{(s)})$, and let $\tilde{w}_{i}^{(s)}$ be the Pareto-smoothed weights (Vehtari et al., 2017). Then the PSIS-LOO estimate is
$\widehat{\mathrm{elpd}}_{\mathrm{PSIS-LOO}}\;=\;\sum_{i=1}^{n}\log\!\Bigl{[}\tfrac{1}{S}\sum_{s=1}^{S}\tilde{w}_{i}^{(s)}\,p(y_{i}\mid\theta^{(s)})\Bigr{]}.$
Instead of refitting the model $n$ times (cost $O(n)$), PSIS-LOO only needs a single full-data fit plus simple weight calculations. The fitted Pareto shape parameters $k_{i}$ are used to check reliability of this method of ELPD computation, and the common rule is that if 20% of the $k_{i}$ exceed $0.7$, the estimate should be considered unreliable.
3 RefineStat
Figure 1: The workflow of RefineStat. (1) Data and prompt are provided to the language model, which generates a probabilistic program. (2) Constrained semantic decoding enforces syntactic and semantic validity of the generated program. (3) A Bayesian reliability check diagnoses convergence, divergences, and predictive validity. If failures are detected, the model is refined by backtracking and resampling priors or likelihoods. (4) Upon passing checks, the final reliable probabilistic program is obtained.
We present RefineStat’s two main ideas. First, we prune the possible search space of probabilistic programs by enforcing semantic validity during generation, mapping validation rules to nodes in the partial parse tree and resampling problematic program fragments when constraints are violated. Second, we implement diagnostic-aware refinement, systematically resampling components of statistically unsound models to satisfy established Bayesian workflow guidelines. This integrated approach aims to achieve semantic correctness, statistical reliability, and strong predictive performance.
3.1 Problem Statement
Let $\mathcal{D}$ denote the dataset for a given statistical modeling task. The objective of RefineStat is to construct a statistical model $M$ in a probabilistic programming language that provides accurate predictive performance while quantifying uncertainty in a fully Bayesian manner.
Our approach to finding a model $M$ that explains the data will follow the standard Bayesian workflow (Gelman et al., 2020) [^1] The Bayesian workflow (Gelman et al., 2020) consists of several stages: (1) Specify model by defining likelihood and priors (in our case using an LLM). (2) Perform posterior inference, (3) Perform Posterior predictive checks. (4) Perform convergence and sampler diagnostics. (5) Compute $\widehat{\mathrm{elpd}}$(M) only if diagnostics pass. (6) Rank models by $\widehat{\mathrm{elpd}}$ (with uncertainty). The key challenge to finding such a model $M$ is to automatically compare various candidate models that an LLM produces. To automate this task, we will use a battery of diagnostics from statistical literature (Section 2.2), computed during the posterior inference in the standard Bayesian workflow (Gelman et al., 2020).
Definition 2 (Bayesian Workflow Reliability score)
Let $\mathcal{M}$ be the set of candidate models, and fix thresholds $\alpha_{R},\beta_{\mathrm{bulk}},\beta_{\mathrm{tail}},\gamma,L_{\textit{cd}}{},\epsilon$. For each $M\in\mathcal{M}$, define seven binary indicators $s_{j}(M)\in\{0,1\}$ corresponding to (see Section 2.2 for the metrics definitions):
$1.\;\max_{\phi}\widehat{R}_{\phi}(M)\leq\alpha_{R},$ $2.\;\mathrm{BFMI}(M)>\gamma,$ $3.\;\min_{\phi}\mathrm{ESS}_{\mathrm{bulk},\phi}(M)\geq\beta_{\mathrm{bulk}}$
$4.\;\mathrm{divergences}(M)=0.,$ $5.\;\min_{\phi}\mathrm{ESS}_{\mathrm{tail},\phi}(M)\geq\beta_{\mathrm{tail}},$ $6.\;\widehat{\mathrm{elpd}}(M)\text{ is finite},$
$7.\;\tfrac{1}{n}\sum_{i=1}^{n}\mathbb{I}[\widehat{k}_{i}(M)\leq L_{\textit{cd}}{}]\geq 1-\epsilon,$
Then the reliability score is $\mathcal{B}(M)=\sum_{j=1}^{7}s_{j}(M)\;\in\{0,1,\dots,7\}.$
These diagnostics can be extracted directly from the MCMC engines (e.g. Stan Carpenter et al. (2017a), PyMC Salvatier et al. (2016)). Although $\widehat{\mathrm{elpd}}$ provides a principled Bayesian measure of out-of-sample predictive accuracy, its Monte Carlo estimate can be unreliable if the sampler has not fully converged or if the importance weights are unstable (Gelman et al., 1995). To mitigate these risks, we consider elpd estimates for models that satisfy standard convergence thresholds, thus ensuring that predictive comparisons rest on reliable posterior samples.
We require each model to pass most of these seven checks: if any check fails, the corresponding $s_{j}(M)$ is zero, and the total score reflects how many diagnostics remain satisfactory. In particular, if $\widehat{\mathrm{elpd}}$(M) cannot be computed (or is infinite), then $s_{\mathrm{ELPD}}(M)=0$. A higher score means that more diagnostics have been verified, and the resulting $\widehat{\mathrm{elpd}}$ estimate can be trusted with greater confidence. We consider a model sufficiently reliable once its score exceeds a cutoff $\zeta$. We will use $\zeta=5$, to accommodate for marginal diagnostic failures while maintaining confidence in $\widehat{\mathrm{elpd}}$.
Definition 3 (Valid model space)
Let $\mathcal{M}$ be the set of candidate models, then
$\mathcal{M}_{\mathrm{valid}}=\bigl{\{}\,M\in\mathcal{M}:\mathcal{B}(M)\geq\zeta\,\bigr{\}}.$
Definition 4 (RefineStat Objective)
We finally define our objective to identify the model that attains the highest ELPD-LOO estimate within the space of models $\mathcal{M}_{\mathrm{valid}}$:
$\framebox{$M^{*}=\arg\max_{M\in\mathcal{M}_{\mathrm{valid}}}\widehat{\mathrm{elpd}}(M),$}$
3.2 Semantically-Constrained Probabilistic Program Generation
Figure 2: A constrained semantic decoding iteration in RefineStat
We formalize the generation of semantically valid probabilistic programs through iterative constrained sampling. Let $G=(\mathcal{N},\mathcal{T},\mathcal{P},S_{0})$ be a context-free grammar with nonterminal symbols $\mathcal{N}$, terminal symbols $\mathcal{T}$, production rules $\mathcal{P}$, and start symbol $S_{0}$. For a partial program $c\in L_{p}(G)$ with parse tree $\kappa$, we define validation functions that operate on parse tree nodes $n\in N(\kappa)$ within program context $\pi\in\Pi$:
$\Phi:N(\kappa)\times\Pi\mapsto\{0,1\}$
These functions are constructed as conjunctions of individual correctness checks:
$\Phi(n,\pi)=\phi^{1}(n,\pi)\land\phi^{2}(n,\pi)\land\cdots\land\phi^{m}(n,\pi)$
Here, the parse tree node $n$ is a program fragment which is semantically valid when all validation functions yield true.
Validity Predicates for Probabilistic Program Fragments. To ensure the correctness of synthesized probabilistic program fragments, we define three essential validation predicates. Each fragment $s\in S_{N}=\{\,s\in\mathcal{S}\mid s\text{ is derivable from nonterminal }N\in\mathcal{N}\,\}$ corresponds to a single statement in the program. Let $\mathcal{F}(s)$ denote the set of all probability distribution functions invoked within fragment $s$. The validation predicates are:
  1. 1.
    Parse-ability: $\phi_{1}(s,\Pi)=1$ if the fragment conforms to the grammar $G$.
  2. 2.
    Distribution validity: $\phi_{2}(s,\Pi)=\prod_{f\in\mathcal{F}(s)}\mathbf{1}\{f\in\mathcal{M}\}$ verifies that each probabilistic operation $f$ exists in the available library $\mathcal{M}$ of PPL.
  3. 3.
    Parameter validity: $\phi_{3}(s,\Pi)=\prod_{f\in\mathcal{F}(s)}\mathbf{1}\{P(f)\subseteq P_{\mathrm{acc}}(f)\}$ confirms that operation parameters $P(f)$ adhere to the accepted specifications $P_{\mathrm{acc}}(f)$, essential for maintaining probabilistic semantics. i.e. we ensure that the provided parameters for any distribution are correct according to the specification of the distribution. For instance, as shown in the Figure 2, a use of parameter "std" was deemed invalid for the Normal distribution and it was resampled correctly as "sigma".
The complete validation predicate $\Phi(s,\Pi)=\phi_{1}(s,\Pi)\land\phi_{2}(s,\Pi)\land\phi_{3}(s,\Pi)$ ensures that generated fragments satisfy all requirements of the probabilistic programming language. Our generation approach leverages these semantics by maintaining a global symbol table $\Pi:\mathcal{A}\to\mathcal{M}$ mapping each alias $a\in\mathcal{A}$ to its module or namespace $m\in\mathcal{M}$.
Generation proceeds via local rejection sampling on $S_{N}$: we repeatedly sample $s\sim S_{N}$ until finding $s^{*}$ where $\Phi(s^{*},\Pi)=1$. This iterative process continues until a termination fragment is generated, ensuring every component in the final probabilistic program satisfies all semantic constraints. Note that our local rejection sampling is token-efficient and we backtrack and precisely resample the tokens that correspond to the violation of the constraints. The approach is particularly effective for languages like PyMC, where maintaining consistent probabilistic variable scopes and dependencies is critical. Combining syntactic constraints with semantic validation enables efficient exploration of the program space while ensuring the probabilistic soundness of generated models.
3.3 Program Validation and Guided Resampling
Algorithm 1 RefineStat Synthesis via $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$ with Constrained Decoding
[…]
3:while $r<R_{\max}$ and $\lvert\mathcal{V}\rvert<\beta$ do
4:  $\mathrm{Prog}\leftarrow\mathcal{D}\|\mathcal{P}\|\mathcal{L}$
5:  if $\neg\Phi(\mathrm{Prog})$ then
6:   $r\leftarrow r+1$; continue
7:  compute diagnostics $d_{1},...,d_{L}$ on $\mathrm{Prog}$
8:  $P_{\rm pass}\leftarrow\bigl{|}\{j:d_{j}\geq\tau_{j}\}\bigr{|}$
9:  if $P_{\rm pass}\geq K$ then
10:   $\mathcal{V}\leftarrow\mathcal{V}\cup\{\mathrm{Prog}\}$; continue
11:  if $\ell<\alpha$ then $\triangleright$ likelihood resampling
12:   $\mathcal{L}\leftarrow L_{\textit{cd}}{}(\mathcal{D}\|\mathcal{P})$
13:   $\ell\leftarrow\ell+1$
[…]
15:   $\mathcal{P}\leftarrow L_{\textit{cd}}{}(\mathcal{D})$
16:  $r\leftarrow r+1$
17:return $\arg\max_{M\in\mathcal{V}}\mathrm{elpd}(M)$
Building on the validation predicate $\Phi$, we now formalize our constrained generation and refinement methodology. We introduce the constrained-decoding operator $\mathcal{L}_{\mathrm{CD}}:\mathcal{C}\to\mathcal{B}$, which implements our validation-guided sampling. Here, a statement is an individual syntactic unit, and a code block $\mathcal{B}$ is a (possibly multi-statement) sequence of such statements. For any context $\mathcal{C}$, this operator returns a new code block $\mathcal{B}$ satisfying $\Phi(\mathcal{C}\|\mathcal{B})=1$, where $\|$ denotes sequential concatenation of blocks; concretely, for two blocks $A$ and $B$, $A\|B$ is the program text formed by appending all statements of $B$ immediately after those of $A$.
In particular, $\mathcal{D}\,\|\,\mathcal{P}\,\|\,\mathcal{L}$ denotes the full probabilistic program consisting first of the data block $\mathcal{D}$, then the prior block $\mathcal{P}$, and finally the likelihood block $\mathcal{L}$. During refinement we perform resampling via:
$\mathcal{L}\;\leftarrow\;L_{\textit{cd}}{}(\mathcal{D}\|\mathcal{P})\quad\text{(likelihood resampling)},$
$\mathcal{P}\;\leftarrow\;L_{\textit{cd}}{}(\mathcal{D})\quad\text{(prior resampling)},$
guaranteeing that each replaced block remains semantically valid.
Algorithm 1 synthesizes programs in two phases. First, it checks semantic correctness via $\Phi$, ensuring parseability, distribution validity, and parameter consistency. Second, it evaluates Bayesian diagnostics $d_{1},\dots,d_{L}$ on the full program $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$; at least $K$ thresholds $\{\tau_{j}\}$ must be met to accept a candidate. If diagnostics fail, we perform one of two resampling steps: (i) likelihood resampling replaces $\mathcal{L}$ under the data–prior context, addressing convergence or sampler-health issues; (ii) prior resampling replaces $\mathcal{P}$ under the data context, correcting prior-specification errors. We iterate until we collect $\beta$ valid programs or exhaust the rejection budget $R_{\max}$, and return the program maximizing $\widehat{\mathrm{elpd}}$.
4 Experimental Methodology
We use PyMC, a Python probabilistic programming library, to perform inference. We provide the same initial prompt while performing unconstrained generation and using RefineStat. We prompt the model by providing it with the dataset, the necessary library and the text query. The exact format of the prompt is provided in the Appendix C. To demonstrate RefineStat’s superiority, we allocate the Baseline a larger token budget than RefineStat, yet RefineStat still outperforms it despite using fewer tokens. Section 5.2 provides a detailed account of the experimental setup, including token allocation.
As defined in Definition 2 (Section 3.1), we assess model reliability using standard Bayesian diagnostics Vehtari et al. (2021, 2017); Gelman et al. (1995). We set the convergence threshold to $\alpha_{R}=1.05$ for split-$\widehat{R}$, allow $\mathrm{ESS}_{\mathrm{bulk}}\geq\beta_{\mathrm{bulk}}=400$, and adopt a relaxed cutoff $\beta_{\mathrm{tail}}=100$ for $\mathrm{ESS}_{\mathrm{tail}}$ to accommodate lower sampling efficiency in the tails. For leave-one-out validation, $\widehat{\mathrm{elpd}}(M)$ must be finite, with at least $1-\epsilon=0.8$ of data points having Pareto shape values below $L_{\mathrm{cd}}=0.7$. These thresholds are used consistently when computing the reliability score across all models. We use Standard unconstrained generation as our baseline. Further, based on preliminary experiments we have chosen $\beta$ to be 4, $\alpha$ to be 2, and $R_{\max}$ as 100 for all experimental purposes.
Datasets. We use five benchmark datasets from the Stan PosteriorDB Magnusson et al. (2024), mirroring their selection in prior research on automated statistical model generation Li et al. (2024):
  • Eight Schools Rubin (1981): This dataset originates from a study commissioned by the Educational Testing Service, which examines the effects of coaching programs on test performance.
  • Dugongs Unit (b): This dataset provides measurements on the ages and lengths of 27 dugongs.
  • Surgical Unit (a): This dataset comprises records on the number of cardiac surgeries performed on infants, along with the associated failure rates.
  • Peregrine M Kery (2011): This dataset tracks the breeding trajectory of the peregrine falcon population in the French Jura region from 1964 to 2003.
  • GP: This dataset contains simulated observations generated from a Poisson Gaussian Process.
Models. We experiment with a range of state-of-the-art LLMs, spanning multiple parameter scales including Qwen2.5 (code-specific) Hui et al. (2024), models from the Llama series, DeepSeek, and Google’s CodeGemma. We have used a total of four models including, Llama3-8B Grattafiori et al. (2024), CodeGemma-7B Team et al. (2024), Qwen2.5-Coder-7B Hui et al. (2024), and DeepSeek-R1-Distill-Qwen-7B (hereafter we refer to it as “DQ-7B”) Guo et al. (2025).
Experimental Setup. We run experiments on a 48-core Intel Xeon Silver 4214R CPU with 2 NVidia RTX A5000 GPUs. RefineStat is implemented using PyTorch Paszke et al. (2019), and Itergen library Ugare et al. (2024a) for refining the parser-guided LLM generation infrastructure. We run all experiments for 10 seeds to reduce result randomness, and use a temperature range of 0.2 to 0.4.
5 Experimental Results
In this section, we demonstrate how RefineStat enhances probabilistic programming model synthesis. We conduct a comprehensive evaluation on diverse datasets, comparing RefineStat against both an unconstrained baseline and leading language models (e.g., OpenAI o3) across a suite of diagnostic and performance metrics. Section 5.4 presents an ablation study that quantifies the contribution of each algorithmic component, metadata and contrasts RefineStat’s token consumption with that of the baseline model. Finally, we demonstrate that our method does not simply recall training data.
5.1 Improvement in Run Rate over Unconstrained and Syntax-driven Generation
To demonstrate the effectiveness of our method, we used identical prompts across our framework, the Standard baseline, Syncode Ugare et al. (2024c), and RefineStat. We compiled and ran each generated program to obtain the Run rates across different methods.
Table 1: Run rates for Standard, SynCode, and RefineStat by temperature
Temp. Standard SynCode RefineStat
0.2 0.10 0.21 0.45
0.3 0.11 0.21 0.50
0.4 0.11 0.21 0.50
Table 1 presents Run rates across temperature settings. Run rate is the fraction of programs that successfully produce the samples from the posterior distribution. The problems that do not run successfully include those with runtime errors such as (1) numerical errors, e.g., inf/nan, (2) sampling issues due to unlikely prior parameterization, (3) other sampling warnings, e.g., failed to initialize chains, (4) static compilation issues. RefineStat achieves success rates approximately 40 percentage points higher than the Standard baseline and 30 points higher than Syncode, demonstrating that our validation-guided approach substantially enhances code generation reliability by mitigating both syntactic and semantic error sources.These results show that RefineStat significantly enhances code generation reliability. Further, we categorized the failures by their root causes across different methods:
  • The Standard baseline exhibited frequent syntax errors (e.g., unmatched delimiters, missing imports) and invalid API calls.
  • Syncode eliminated many basic syntactic mistakes but still suffered semantic errors, such as incorrect distribution parameter names, type mismatches, referring to deprecated API functions (e.g., calling pm.sample_prior from an earlier PyMC release), and inventing non-existent methods like pm.random_coefs.
  • RefineStat, in contrast, often produced models whose samplers failed to explore the correct posterior modes, leading to chains stuck in low-density regions or divergent transitions, failures stemming from the model definitions rather than our decoding framework; RefineStat reduced both syntactic and semantic errors and avoided sampler pathologies by enforcing grammar and parameter validity during decoding.
5.2 Comparison of Generated Program Quality to Unconstrained Baseline
To evaluate our framework’s ability to produce higher quality probabilistic programs in terms of semantic correctness and diagnostic robustness, we ran two experimental conditions using the same prompt: the base language model alone, and RefineStat. Upon experimentation with RefineStat, we observed that on an average it consume almost twice the number of tokens consumed by the Baseline. Thus, to provide advantage to baseline models, we select the best program in terms of Bayesian Reliability Score, and $\widehat{\mathrm{elpd}}$ value among five different seeds for each dataset for comparison to a single run of RefineStat. We repeat this process five times for different seeds, to obtain the mean and standard deviation for all the metrics computed using these generated programs.
We then compare each generated program against the Bayesian Workflow Reliability score and the full suite of seven diagnostic criteria introduced earlier. To conserve space, we report five representative metrics drawn from these diagnostics. In our tables, the symbol ✗ indicates that no valid program was produced—i.e., the method failed to explore the search space sufficiently to yield a correct result.
DQ-7B failed on every dataset under the Standard Baseline, whereas it succeeded on all datasets when augmented with RefineStat. Except for one instance, RefineStat matches or exceeds the Standard Baseline in terms of the Bayesian Workflow Reliability score, and in some cases achieves up to twice the reliability. Moreover, RefineStat delivers substantially better performance on individual diagnostics, particularly divergences. For example, on the Surgical dataset with Meta-Llama, the Standard Baseline produces over 1000 divergences, while RefineStat produces none.
We observed that in cases where the Standard baseline attains a higher ELPD-LOO than RefineStat, closer inspection of the diagnostics exposes unreliable sampling. For instance, on the Meta-Llama GP task the Standard model achieves a superior ELPD score, but exhibits split-$\widehat{R}=4.13\gg 1$ and a low reliability score (3), indicating severe convergence issues. In contrast, RefineStat may report a marginally lower ELPD yet maintains $\widehat{R}\approx 1$ and a higher reliability score, reflecting trustworthy posterior estimates. Furthermore, RefineStat delivers markedly lower variability in key diagnostics such as $\widehat{R}$ and the number of divergent transitions demonstrating its consistency and robustness.
Table 2: Comparison of Diagnostic Scores and ELPD-LOO for Standard vs. RefineStat
Dataset Model Variant Reliab. Score $\uparrow$ $\widehat{R} \downarrow$ ESS Bulk $\uparrow$ Divergences $\downarrow$ Pareto $k$ $\downarrow$ ELPD LOO $\uparrow$
Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std
8 Schools Meta-LLama-3-8B Standard 7.00 0.00 1.00 0.00 2261.00 0.00 0.00 0.00 0.00 0.00 -31.70 0.00
RefineStat 7.00 0.00 1.00 0.00 2303.00 768.76 0.00 0.00 0.00 0.05 -31.77 0.61
CodeGemma-7B Standard
RefineStat 7.00 0.49 1.00 0.00 2573.00 527.68 0.00 1.97 0.00 0.06 -31.46 1.84
Qwen-Coder-7B Standard 3.80 1.30 1.02 0.01 223.25 80.43 92.75 31.38 0.22 0.21 -31.31 0.68
RefineStat 5.00 0.45 1.01 0.01 256.50 883.83 34.50 86.89 0.00 0.05 -30.80 0.05
DQ-7B Standard
RefineStat 5.00 0.98 1.02 0.02 219.00 2176.00 102.00 75.36 0.00 0.00 -30.68 0.11
Dugongs Meta-LLama-3-8B Standard 7.00 0.00 1.00 0.00 2167.67 884.79 0.00 0.00 0.01 0.02 -7.66 30.39
RefineStat 7.00 0.00 1.00 0.00 1696.00 284.33 0.00 0.00 0.00 0.02 8.42 24.51
CodeGemma-7B Standard 5.70 2.30 1.04 0.06 1527.33 1325.54 285.67 494.79 0.00 0.00 3.90 7.71
RefineStat 7.00 0.00 1.00 0.00 1908.00 2066.23 0.00 0.00 0.04 0.02 8.07 15.42
Qwen-Coder-7B Standard 7.00 0.00 1.00 0.01 1788.67 123.43 0.00 0.00 0.04 0.00 8.15 0.29
RefineStat 7.00 0.00 1.00 0.00 1683.00 148.16 0.00 0.00 0.04 0.02 8.29 0.05
DQ-7B Standard
RefineStat 7.00 0.00 1.00 0.00 2376.00 149.61 0.00 0.00 0.00 0.03 8.35 0.06
Peregrine Meta-LLama-3-8B Standard 6.00 0.00 1.00 0.00 3774.00 0.00 7.00 0.00 0.00 0.00 -184.96 0.00
RefineStat 7.00 0.00 1.00 0.00 3574.00 428.26 0.00 0.00 0.00 0.00 -173.00 4.91
CodeGemma-7B Standard 7.00 0.00 1.00 0.00 4261.00 0.00 0.00 0.00 0.00 0.00 -172.91 0.00
RefineStat 6.50 0.53 1.00 0.00 2930.00 1343.79 0.50 0.53 0.00 0.00 -129.93 3.91
Qwen-Coder-7B Standard 7.00 0.00 1.00 0.00 4238.00 0.00 0.00 0.00 0.00 0.00 -173.11 0.00
RefineStat 7.00 0.00 1.00 0.00 4679.00 88.73 0.00 0.00 0.00 0.00 -172.98 0.12
DQ-7B Standard
RefineStat 6.50 0.53 1.01 0.01 963.50 436.70 25.50 27.26 0.00 0.00 -114.29 2.76
Surgical Meta-LLama-3-8B Standard 5.50 1.30 1.01 0.01 1159.75 1202.11 1066.50 1267.61 0.65 0.44 -31.59 20.02
RefineStat 6.00 0.77 1.00 0.00 579.00 1272.12 0.00 5.75 0.00 0.36 -46.73 36.63
CodeGemma-7B Standard 5.50 0.70 1.00 0.00 1230.00 503.46 6.00 5.66 0.46 0.65 -42.02 5.75
RefineStat 6.00 0.49 1.00 0.00 2026.00 450.46 0.00 0.49 0.00 0.12 -45.55 381.60
Qwen-Coder-7B Standard 6.30 1.50 1.01 0.02 1332.75 784.13 37.50 75.00 0.23 0.46 -44.48 4.35
RefineStat 7.00 0.41 1.00 0.00 1642.00 1350.46 0.00 0.00 0.00 0.38 -46.55 2.79
DQ-7B Standard
RefineStat 7.00 0.50 1.00 0.01 2800.00 1182.68 0.00 0.00 0.00 0.37 -46.51 3.49
GP Meta-LLama-3-8B Standard 3.00 0.00 4.13 0.00 4.00 0.00 1634.00 0.00 0.00 0.00 -21.61 0.00
RefineStat 6.00 0.49 1.00 0.00 1710.00 668.57 13.00 12.39 0.09 0.08 -152.30 139.07
CodeGemma-7B Standard 7.00 0.00 1.00 0.00 6034.00 0.00 0.00 0.00 0.18 0.00 -154.42 0.00
RefineStat 7.00 0.98 1.00 0.00 1752.00 1004.88 0.00 2.46 0.00 0.49 -22.76 126.08
Qwen-Coder-7B Standard
RefineStat
DQ-7B Standard
RefineStat 6.50 0.53 1.01 0.00 816.50 89.27 30.50 32.61 0.00 0.00 -23.39 1.14
5.3 Comparison of Generated Program Quality to BoxLM
Table 3: Comparison with BoxLM Li et al. (2024), RefineStat using DQ-7B, and Expert values
Dataset Expert RefineStat BoxLM OpenAI-o3
w/ DQ-7B w/ GPT-4
Eight schools -30.70 -30.68 -30.42 -30.74
Dugongs 22.43 8.35 23.40 22.83
Peregrine -112.60 -114.29 -173.11 -133.29
Surgical -39.73 -46.51 -38.03 -38.73
GP -26.53 -23.39 -34.95
In this section, we compare the performance of our framework using the DQ-7B model (averaged over five runs) against three baselines: the Expert programs from PosteriorDB Magnusson et al. (2024), the BoxLM system introduced by Li et al. (2024), and programs generated by OpenAI o3. Since the code for BoxLM is not publicly available, we rely on the reported numbers from their paper for comparison. Additionally, because the specific dataset we use for the GP task is not included in BoxLM’s evaluation suite, we omit their result for that dataset in Table 3.
As shown in Table 3, our framework consistently outperforms both BoxLM and OpenAI o3 on the Peregrine dataset, and surpasses OpenAI o3 on the GP task, while matching the performance of expert-written programs in most cases. Across the remaining datasets, our approach performs comparably to other baselines, with the exception of Dugongs, where performance is slightly lower. These results demonstrate that RefineStat achieves performance comparable to, and in several cases better than, large language models like OpenAI o3 and multi-agent frameworks such as BoxLM.
5.4 Ablation Study
Effectiveness of Semantic Validation Components. To evaluate the contribution of each validity predicate within RefineStat, we perform an ablation study by systematically disabling one component at a time and measuring the resulting compilation rate. This analysis is conducted across all models with ten different random seeds to ensure robustness. Notably, removing all validation predicate reduces the system to the behavior of SynCode.
Table 4 shows that parameter validity emerges as the most critical component, its removal results in a substantial drop of 14.5 percentage points in compilation success. Nonetheless, all checks contribute meaningfully to overall performance: omitting distribution validity or parse-ability leads to consistent declines in compilation rates (−9% and −5.5%, respectively), underscoring the complementary role these validations play in ensuring soundness of generated programs.
Table 4: Ablation Study: Impact of Semantic Validation Checks on Run rate
[…]
1 RefineStat (all components) 50.0% -
2 w/o Parameter validity 35.5% -14.5%
3 w/o Distribution validity 26.5% -9%
4 w/o Parse-ability 21.0% -5.5%
5 w/o grammar-guided generation 11.0% -10.0 %
Token Efficiency Analysis. To evaluate the computational cost associated with our framework, we measure the number of tokens consumed in generating the best-performing program under Itergen, Baseline (Unconstrained generation), RefineStat without Refinement Loop (RefineStat w/o RL), and RefineStat using Meta-LLama-3-8B. The cumulative token usage across all ten runs is recorded, and this process is repeated across five independent batches (each using a new set of seeds), from which we report the mean and standard deviation.
In contrast, RefineStat is run five times using five independent seeds, and we track the total number of tokens consumed throughout the entire refinement process in each run.
The Table 5 presents the results across models and datasets, with the final column (“Token Ratio”) reporting the ratio of token usage by RefineStat relative to the baseline. On average, RefineStat consumes twice the number of tokens consumed by Baseline. While this reflects the added cost of our refinement mechanism, the overhead varies across settings. For instance, in some cases such as Dugongs, RefineStat uses fewer tokens than the baseline due to early convergence. Conversely, high multipliers (e.g., Eight Schools, GP) reflect continued refinement due to unmet stopping conditions, even if the generated program is already of high quality.
Table 5: Comparison of Itergen, Baseline, and RefineStat Variants with Multipliers.
Dataset Itergen Baseline RefineStat w/o RL RefineStat Token
Mean Std Mean Std Mean Std Mean Std Ratio
Eight Schools $98.8$ $5.3$ $606.8$ $386.3$ $147.6$ $36.7$ $1503.0$ $409.5$ $2.5$
Dugongs $209.3$ $225.9$ $747.4$ $495.0$ $294.4$ $86.3$ $419.0$ $87.1$ $0.6$
GP $131.3$ $8.9$ $663.6$ $410.9$ $155.0$ $21.2$ $1660.0$ $253.5$ $2.5$
Peregrine $132.5$ $21.4$ $884.0$ $538.9$ $180.8$ $103.6$ $1740.0$ $418.0$ $2.0$
Surgical $97.3$ $4.3$ $624.4$ $260.6$ $113.2$ $9.2$ $1275.0$ $1348.1$ $2.0$
Average Token Ratio 1.9x
5.5 Memorization Effect
While canonical textbook datasets such as Eight Schools and Radon are valuable for reproducible benchmarking and comparison to prior work (e.g., Li et al. (2024); PyMC3 implementations), there remains a concern that a model could rely on memorization rather than genuine synthesis when such datasets are well known. Prior work, including BoxLM, has also included dataset descriptions directly in prompts (see Figure 1 in Li et al., 2024).
To assess the extent to which RefineStat depends on memorization, we performed two controlled prompt modifications on all datasets using Meta-Llama 3-8B:
Anonymized Prompt (RefineStat-AP): All metadata and dataset names were removed, leaving only the raw dataset.
Syntactic Obfuscation (RefineStat-SO): All numerical values were transformed into exponential notation (e.g., 3.28e2 instead of 328) to prevent exact string matches with any potential training data.
Anonymized Prompt and Syntactic Obfuscation (RefineStat-AP-SO): Combined variant using both Anonymized Prompt, and Syntactic Obfuscation.
As shown in Table 6, both variants match the original RefineStat on reliability, convergence, and predictive metrics. The table also reports a combined variant using both modifications, which performs comparably. This consistency suggests that RefineStat’s effectiveness stems from learning from the provided data rather than memorization.
Table 6: Comparison of Diagnostic Scores and ELPD-LOO for RefineStat variants.
Dataset Variant Reliab. Score $\uparrow$ $\widehat{R} \downarrow$ ESS Bulk $\uparrow$ Divergences $\downarrow$ Pareto $k$ $\downarrow$ ELPD LOO $\uparrow$
Mean Std Mean Std Mean Std Mean Std Mean Std Mean Std
Dugongs RefineStat 7.00 0.00 1.00 0.00 1696.00 284.33 0.00 0.00 0.00 0.02 8.42 24.51
RefineStat-AP 7.00 0.00 1.00 0.00 1073.00 317.84 0.00 0.00 0.04 0.00 -0.17 4.24
RefineStat-SO 7.00 0.00 1.00 0.00 1753.50 68.95 0.00 0.00 0.04 0.00 8.31 0.09
RefineStat-AP-SO 7.00 0.00 1.00 0.00 2257.00 731.23 0.00 0.00 0.00 0.00 1.79 7.09
Eight Schools RefineStat 7.00 0.00 1.00 0.00 2303.00 768.76 0.00 0.00 0.00 0.05 -31.77 0.61
RefineStat-AP 7.00 0.00 1.00 0.00 2926.00 541.38 0.00 0.00 0.00 0.00 -31.62 0.79
RefineStat-SO 7.00 0.00 1.00 0.00 2470.50 32.61 0.00 0.00 0.06 0.06 -31.61 0.01
RefineStat-AP-SO 7.00 0.00 1.00 0.00 2187.50 252.83 0.00 0.00 0.00 0.00 -31.60 0.04
Peregrine RefineStat 7.00 0.00 1.00 0.00 3574.00 428.26 0.00 0.00 0.00 0.00 -173.00 4.91
RefineStat-AP 7.00 0.00 1.00 0.00 4057.00 811.85 0.00 0.00 0.00 0.00 -173.14 65.91
RefineStat-SO 7.00 0.00 1.00 0.00 1812.50 24.05 0.00 0.00 0.00 0.00 -132.88 8.20
RefineStat-AP-SO 6.00 0.00 1.00 0.00 1812.50 24.05 0.00 0.00 0.00 0.00 -140.88 8.20
GP RefineStat 6.00 0.49 1.00 0.00 1710.00 668.57 13.00 12.39 0.09 0.08 -152.30 139.07
RefineStat-AP 7.00 0.00 1.00 0.00 2283.00 519.55 0.00 0.00 0.00 0.04 -21.24 2.58
RefineStat-SO 7.00 0.00 1.00 0.00 1135.00 0.00 0.00 0.00 0.00 0.00 -24.99 0.00
RefineStat-AP-SO 6.50 0.53 1.00 0.00 1140.00 251.23 69.00 73.76 0.00 0.00 -23.01 0.47
Surgical RefineStat 6.00 0.77 1.00 0.00 579.00 1272.12 0.00 5.75 0.00 0.36 -46.73 36.63
RefineStat-AP 7.00 0.49 1.01 0.00 640.00 635.48 0.00 0.00 0.00 0.25 -46.71 96.47
RefineStat-SO 7.00 0.00 1.01 0.01 1311.00 742.99 0.00 0.00 0.04 0.04 -65.22 19.85
RefineStat-AP-SO 7.00 0.00 1.01 0.01 1139.00 638.22 0.00 0.00 0.04 0.04 -69.27 24.20
6 Related Works
Researchers have developed several probabilistic techniques for discovering model structures from observed data, including Bayesian networks (Mansinghka et al., 2006; Lowd and Domingos, 2012), matrix-composition models (Grosse et al., 2012), Markov networks (Gogate et al., 2010), and deep probabilistic models (Gens and Pedro, 2013). Probabilistic programming languages (PPLs) (Salvatier et al., 2016; Phan et al., 2019; Carpenter et al., 2017a) provide ways to flexibly represent these models as programs and support generic inference methods. While these approaches have shown promise, they often require both significant domain expertise and specialized API and/or programming language expertise to implement effectively. Given the recent advances in large language models’ (LLMs) code generation capabilities, synthesizing probabilistic programs using LLMs offers a potentially more accessible and flexible approach to automated model discovery. Therefore, our work RefineStat addresses this opportunity by advancing the SOTA in LLM-based probabilistic program synthesis.
Probabilistic Program Synthesis: There have been many approaches for deterministic program synthesis, which is recently been dominated by LLM-based approaches. Several works (Nori et al., 2015; Saad et al., 2019; Gerasimou et al., 2015; Češka et al., 2019; Ellis et al., 2015) have attempted to synthesize probabilistic programs using classical machine learning or symbolic methods. PSKETCH Nori et al. (2015) applies the notion of sketching for probabilistic program synthesis. Gerasimou et al. (2015) uses a multiobjective optimization genetic algorithm for probabilistic model generation. Saad et al. (2019) presents a technique for automatically constructing probabilistic programs for data analysis, interpretation, and prediction.
Most recently, Li et al. (2024) uses LLM for probabilistic program synthesis. The key idea is to use two interacting LLMs: a generator LLM and a critic LLM that improve the synthesized program iteratively. The paper shows that with instances of GPT4 as the generator and critic (closed LLM with likely >>T weights) can find reasonable probabilistic programs from data. As our evaluation shows, RefineStat significantly improves the ability to find programs that fits the data (recall Table 3) and in contrast to Li et al. (2024), runs only a single small open LLM (<8<8B weights), demonstrating the positive impact of constrained decoding.
Program Synthesis with Constrained LLM Decoding: Recent advances in program synthesis have enabled constrained decoding approaches where LLMs generate code while adhering to formal language specifications. These constraints can be partially precomputed and enforced more efficiently for regular (Deutsch et al., 2019; Willard and Louf, 2023; Kuchnik et al., 2023) or context-free (Koo et al., 2024; Ugare et al., 2024b; Dong et al., 2024; Banerjee et al., 2025; Suresh et al., 2025; Firestone et al., 2025) languages, ensuring syntactic correctness. For more dynamic program generation, Poesia et al. (2022) and Ugare et al. (2025) implement error-driven backtracking when the LLM synthesizes invalid code.
[…]
7 Conclusion and Limitations
Conclusion.
The main contribution of our work is to separate the task of generating probabilistic modeling through PPLs as fragments of priors and likelihood and to construct an LLM-based search procedure that automatically discovers the probabilistic program that satisfies the standard reliability metrics in the Bayesian workflow. We believe that our framework can be extended to enforce arbitrary reliability criteria defined by domain experts for reliable generation in other domains that involve domain-specific languages and plan to explore those in future work.
Limitations.
While our framework incorporates key components of the Bayesian workflow, such as convergence diagnostics and predictive performance metrics, it does not currently include prior-predictive or posterior-predictive checks, which often require manual inspection and domain-specific judgment Gelman et al. (1995). As a result, the reliability assessment is based on a subset of available diagnostics, and in some cases, the reported ELPD values may not fully reflect model adequacy. Moreover, our refinement strategy, while effective in practice, does not guarantee convergence to a fully satisfactory or globally optimal program.
Appendix
This appendix provides expanded technical details, evaluations, and examples supplementing the main paper. Below is a structured index:
A. Probabilistic Program Metrics
  • $\circ$
    $\widehat{R}$ Statistic
  • $\circ$
    Effective Sample Size (ESS)
  • $\circ$
    Divergent Transitions
  • $\circ$
    Bayesian Fraction of Missing Information (BFMI)
  • $\circ$
    Pareto Shape Parameter
  • $\circ$
    Expected Log Pointwise Predictive Density (ELPD)
B. Error Analysis
  • $\circ$
    Search Space Limitations
  • $\circ$
    Model Misfit and Sampling Failures
  • $\circ$
    Call-Level Hallucinations
  • $\circ$
    Termination Failures Due to Budget Constraints
C. Prompt Design
  • $\circ$
    Prompt Template
  • $\circ$
    Dataset-Specific Templates:
Appendix A Probabilistic Program Metrics
In our framework for valid statistical model synthesis, we employ several diagnostic metrics from probabilistic programming to ensure model validity. Below, we define each of these metrics formally.
Definition 5 ($\widehat{R}$ Statistic)
The split-$\widehat{R}$ statistic for parameter $\phi$, denoted $\widehat{R}{\phi}$, measures the convergence of Markov chains in MCMC sampling by comparing the between-chain variance to the within-chain variance. Formally:
$\widehat{R}{\phi}=\sqrt{\frac{V}{W}}$ (1)
where $V$ is the variance between chain means and $W$ is the average variance within chains. Values close to 1.0 indicate convergence, while higher values suggest poor mixing of chains.
Definition 6 (Effective Sample Size)
The effective sample size (ESS) measures the equivalent number of independent samples obtained from autocorrelated MCMC draws. For parameter $\phi$, we define:
$\mathrm{ESS}_{\mathrm{bulk},\phi}$ $=\frac{MN}{\tau_{\mathrm{bulk},\phi}}$ (2)
$\mathrm{ESS}_{\mathrm{tail},\phi}$ $=\frac{MN}{\tau_{\mathrm{tail},\phi}}$ (3)
where $M$ is the number of chains, $N$ is the number of draws per chain, and $\tau$ represents the autocorrelation time for bulk or tail estimates respectively. Bulk ESS evaluates sampling efficiency across the central mass of the posterior, while tail ESS focuses on the distribution tails.
Definition 7 (Divergent Transitions)
Divergent transitions, denoted $\mathrm{divergences}(M)$ for model $M$, count the number of leapfrog steps in Hamiltonian Monte Carlo where the numerical approximation of Hamiltonian dynamics breaks down due to extremely high curvature in the posterior geometry. These indicate potential pathological geometries in the posterior distribution that may lead to biased inference.
Definition 8 (Bayesian Fraction of Missing Information)
The Bayesian Fraction of Missing Information, $\mathrm{BFMI}(M)$ for model $M$, is defined as:
$\mathrm{BFMI}(M)=\frac{\mathrm{Var}(\Delta E)}{\mathrm{Var}(E)}$ (4)
where $E$ represents the energy (negative log probability density) and $\Delta E$ is the change in energy between consecutive HMC iterations. Low BFMI values indicate poor exploration of the target distribution.
Definition 9 (Pareto Shape Parameter)
The Pareto shape parameter $\widehat{k}_{i}(M)$ for observation $i$ in model $M$ quantifies the reliability of importance sampling estimates used in PSIS-LOO cross-validation:
[…]
Values $\widehat{k}_{i}<0.5$ indicate reliable estimates, while $\widehat{k}_{i}>0.7$ suggest unstable estimates that may require more robust computational approaches.
Definition 10 (Expected Log Pointwise Predictive Density)
The Expected Log Pointwise Predictive Density under Leave-One-Out cross-validation, $\widehat{\mathrm{elpd}}(M)$ for model $M$, measures the model’s out-of-sample predictive accuracy:
$\widehat{\mathrm{elpd}}(M)=\sum_{i=1}^{n}\log p(y_{i}\mid y_{-i})$ (6)
where $p(y_{i}\mid y_{-i})$ is the predictive density for observation $i$ after fitting the model to all other observations. Higher values indicate better predictive performance.
Appendix B Error Analysis
While RefineStat significantly improves the syntactic correctness and statistical validity of generated probabilistic programs, we analyze the remaining failure cases to better understand the limitations of the framework and identify directions for future improvement.
Search Space Limitations.
Despite the use of resampling based mechanism for semantic validity, the language model occasionally reintroduces previously rejected code fragments. For instance, it may repeatedly generate outdated or invalid syntax such as the use of sd instead of sigma in PyMC model definitions:
mu = pm.Normal(‘‘mu’’, mu=0, sd=10)
This behavior likely stems from the model’s prior exposure to deprecated APIs in its pretraining corpus and reflects the difficulty of escaping local attractors in the search space.
Model Misfit and Sampling Failures.
Despite generating semantically correct code, some models fail during posterior inference due to numerical instabilities inherent in the model specification. A common manifestation of this issue is the PyMC error:
SamplingError: Initial evaluation of model at starting point failed!
This error often arises when certain mathematical operations within the model, such as exponentiation or logarithms, result in undefined or non-finite values. For instance, exponentiating large numbers can lead to overflow, while taking the logarithm of zero or negative numbers is undefined. These numerical issues can cause the log-probability evaluations to return NaN or inf, thereby preventing the sampler from initializing properly.
Call-Level Hallucinations.
The model occasionally hallucinates invalid function names or API calls not present in the target probabilistic programming language. For example:
mu = pm.ExtNormal(‘ex’, mu=0)
Such hallucinations highlight a mismatch between the syntactic plausibility and the executable validity of generated code, reinforcing the need for grounded semantic constraints during decoding.
Termination Failures due to Budget Constraints.
In practice, we impose limits on the maximum number of iterations or generated tokens to maintain tractability. In some instances, these constraints are reached before a valid program is synthesized, resulting in truncated or incomplete code outputs.
SamplingError
Initial evaluation of the model at the starting point failed due to numerical instabilities (overflow/NaNs).
Indentation and commenting failures:
Wrong indentation of python-like function codes; inability to always close string comments.
Appendix C Prompt Design
We use the same prompt across both the baseline, and RefineStat for experimentation purpose. To standardize the prompt across different datasets, we use a template in which the fields {description} and {template_code} are replaced with the dataset-specific description and code snippet, respectively.
Prompt Template Template prompt: # Complete the PyMC model definition within the with pm.Model() as m:’ block below. Your output must define a complete Bayesian model with appropriate priors, likelihood, and then sample the posterior using, pm.sample(1000, tune = 1000, chains = 4, return_inferencedata = True, idata_kwargs = {{"log_likelihood": True}})‘. Do not include any extra commentary or text outside the code. Follow best practices for expert-level Bayesian modeling. # Description: {Description} {Template_Code}
Note: The placeholders {description} and {template_code} are automatically substituted for each dataset. Below are the {description} and {template_code} respectively for each dataset:
C.1 Eight Schools
Description: A hierarchical model for the 8-schools data.
Template Code import pymc as pm import numpy as np import arviz as az import matplotlib.pyplot as plt # Data y = np.array([28, 8, -3, 7, -1, 1, 18, 12]) sigma = np.array([15, 10, 16, 11, 9, 11, 10, 18]) with pm.Model() as m:
C.2 Dugongs
Description: A growth model for dugongs with missing data.
Template Code import pymc as pm import numpy as np import arviz as az import matplotlib.pyplot as plt # Data X = np.array([1, 1.5, 1.5, 1.5, 2.5, 4, 5, 5, 7, 8, 8.5, 9, 9.5, 9.5, 10, 12, 12, 13, 13, 14.5, 15.5, 15.5, 16.5, 17, 22.5, 29, 31.5]) y = np.array([1.8, 1.85, 1.87, 1.77, 2.02, 2.27, 2.15, 2.26, 2.47, 2.19, 2.26, 2.4, 2.39, 2.41, 2.5, 2.32, 2.32, 2.43, 2.47, 2.56, 2.65, 2.47, 2.64, 2.56, 2.7, 2.72, 2.57]) with pm.Model() as m:
C.3 Surgical
Description: The mortality rates in 12 hospitals performing cardiac surgery on babies.
Template Code import pymc as pm import numpy as np import arviz as az import matplotlib.pyplot as plt # Given Data N = 12 # Number of observations n = np.array([47, 148, 119, 810, 211, 196, 148, 215, 207, 97, 256, 360]) r = np.array([0, 18, 8, 46, 8, 13, 9, 31, 14, 8, 29, 24]) with pm.Model() as m:
C.4 GP
Description: Simulated data from a Poisson GP model.
Template Code import pymc as pm import numpy as np import arviz as az import matplotlib.pyplot as plt # Given Data N = 11 # Number of observations x = np.array([-10, -8, -6, -4, -2, 0, 2, 4, 6, 8, 10]) y = np.array([4.75906, 1.59423, 2.99548, 5.27501, 1.66472, 2.24347, 2.8914, 4.08681, 4.60588, 0.802364, 3.92136]) k = np.array([40, 37, 29, 12, 4, 3, 9, 19, 77, 82, 33]) with pm.Model() as m:
C.5 Peregrine
Description: Simulated population counts of peregrines in the French Jura over 9 years
Template Code import pymc as pm import numpy as np import arviz as az import matplotlib.pyplot as plt # Data nyears = 40 # Number of years year = np.array([-0.95, -0.9, -0.85, -0.8, -0.75, -0.7, -0.65, -0.6, -0.55, -0.5, -0.45, -0.4, -0.35, -0.3, -0.25, -0.2, -0.15, -0.1, -0.05, 0, 0.05, 0.1, 0.15, 0.2, 0.25, 0.3, 0.35, 0.4, 0.45, 0.5, 0.55, 0.6, 0.65, 0.7, 0.75, 0.8, 0.85, 0.9, 0.95, 1]) C = np.array([27, 42, 35, 55, 61, 19, 41, 74, 43, 42, 73, 37, 48, 49, 19, 72, 30, 18, 31, 71, 63, 51, 48, 73, 49, 54, 43, 59, 30, 24, 62, 55, 51, 47, 14, 27, 45, 20, 26, 19]) N = np.array([43, 83, 53, 91, 95, 24, 62, 91, 64, 57, 97, 56, 74, 66, 28, 92, 40, 23, 46, 96, 91, 75, 71, 100, 72, 77, 64, 68, 43, 32, 97, 92, 75, 84, 22, 58, 81, 37, 45, 39]) with pm.Model() as m: